# Heeji Yoon1*  Jaewoo Jung1*  Junwan Kim2*Hyungyu Choi3 Heeseong Shin1Sangbeom Lim4Honggyu An1Chaehyun Kim1Jisang Han1 Donghyun Kim4

### v1

@@ Abstract
Multimodal large language models (MLLMs) trained with visual instruction tuning have achieved strong performance across diverse tasks, yet they remain limited in vision-centric tasks such as object counting or spatial reasoning. We attribute this gap to the prevailing text-only supervision paradigm, which provides only indirect guidance for the visual pathway and often leads MLLMs to discard fine-grained visual details during training. In this paper, we present VIsual Representation ALignment (VIRAL), a simple yet effective regularization strategy that aligns the internal visual representations of MLLMs with those of pre-trained vision foundation models (VFMs). By explicitly enforcing this alignment, VIRAL enables the model not only to retain critical visual details from the input vision encoder but also to complement additional visual knowledge from VFMs, thereby enhancing its ability to reason over complex visual inputs. Our experiments demonstrate consistent improvements across all tasks on widely adopted multimodal benchmarks. Furthermore, we conduct comprehensive ablation studies to validate the key design choices underlying our framework. We believe this simple finding opens up an important direction for the effective integration of visual information in training MLLMs.
Figure 1: VIsual Representation ALignment (VIRAL) preserves fine-grained visual attributes for multimodal reasoning. (a) VIRAL introduces an auxiliary regularization objective on the visual pathway to prevent MLLMs from discarding detailed attributes during training. (b) VIRAL, when trained with DINOv2 (Oquab et al., 2023) as the vision foundation model (VFM), consistently produces more accurate visually grounded responses and achieves substantial improvements over standard baselines (Liu et al., 2023) across diverse vision encoders, including CLIP (Radford et al., 2021) and SigLIPv2 (Tschannen et al., 2025).
@@ 1 Introduction
Recent advancements in multimodal large language models (MLLMs) (OpenAI, 2023; Bai et al., 2023a; Team et al., 2023; Chen et al., 2024c), particularly those employing visual instruction tuning techniques such as LLaVA (Liu et al., 2023), have achieved notable success in diverse multimodal tasks. By connecting pretrained large language models (LLMs) (Touvron et al., 2023; Chiang et al., 2023; Chen et al., 2024c; Bai et al., 2025) and vision encoders (Radford et al., 2021; Chen et al., 2024c; Tong et al., 2024a) with a lightweight vision-language projector, visual instruction tuning enables LLMs to understand visual context, demonstrating strong performance across diverse tasks (Chen et al., 2024a; 2025a; Li et al., 2025).
Despite these successes, whether MLLMs truly achieve multimodal understanding remains an open question, as numerous studies report persistent limitations in vision-centric tasks such as object counting and spatial reasoning (Tong et al., 2024b; Qi et al., 2025; Yuksekgonul et al., 2022; Ma et al., 2023). Early approaches largely attribute these shortcomings to the visual encoder or the projector. In response, subsequent works have introduced stronger vision encoders (Lu et al., 2024; Li et al., 2024) and more expressive projectors (Liu et al., 2024; Cha et al., 2024; McKinzie et al., 2024), aiming to supply the language model with richer and more comprehensive visual representations. While they yield measurable improvements, approaches that rely solely on more powerful vision encoders or projectors are inherently constrained in scalability and efficiency.
In this paper, we first revisit the conventional training paradigm of visual instruction tuning. Existing MLLMs are predominantly fine-tuned with a language-modeling objective, updating both the LLM and the vision-language projector while concentrating supervision almost entirely on textual outputs (Li et al., 2024; Bai et al., 2023b; Chen et al., 2024c). As a result, visual tokens receive only indirect, language-mediated supervision despite comprising a substantial fraction of the multimodal input. In effect, the visual pathway remains under-supervised, motivating a central question: Is the prevailing multimodal training setup adequate for capturing and preserving visual information?
@@
Motivated by these findings, we propose VIsual Representation ALignment (VIRAL), a simple yet effective regularization strategy that directly supervises the visual pathway in MLLMs to prevent the model from discarding fine-grained visual attributes provided by the vision encoder during training. Specifically, we align the internal visual representations of the MLLMs with those of the initial vision encoder using an alignment loss based on cosine similarity. In addition, we further find that this alignment signal is much more effective when provided from stronger vision foundation models (VFMs) (Oquab et al., 2023; Kirillov et al., 2023; Yang et al., 2024; Ranzinger et al., 2024). Since VFMs are trained on vision-centric objectives, they provide rich visual representations that complement language supervision. Therefore, aligning the internal visual representations of MLLMs with those of VFMs allows the model to retain important visual details from the input vision encoder while also absorbing additional visual knowledge from VFMs, which in turn enhances its ability to reason over complex visual inputs. Through extensive experiments on widely adopted multimodal benchmarks, we show that VIRAL consistently delivers significant improvements across all tasks.
@@
  • We show that, under the visual instruction tuning paradigm, internal visual representations in MLLMs often lose alignment with the rich features produced by vision encoders, leading to the degradation of fine-grained visual information.
  • We propose a novel regularization strategy that explicitly aligns MLLM visual representations with informative features from pretrained VFMs, thereby preventing the loss of fine-grained attributes and enabling richer multimodal understanding.
  • Through comprehensive experiments on standard multimodal benchmarks, we demonstrate consistent improvements, and further conduct extensive ablation studies to validate the effectiveness of our design choices.
@@ 2 Related Work
Recent studies (Kaduri et al., 2025; Zhang et al., 2025b) have revealed a structured processing hierarchy in MLLMs for vision–language inputs: early layers primarily aggregate global visual context into token-level embeddings, intermediate layers capture fine-grained and spatially localized features, and later layers integrate multimodal information to facilitate response generation.
Within this hierarchy, the middle layers have been shown to be particularly critical for visual understanding. Jiang et al. (2025) decompose these layers into enrichment and refinement phases, demonstrating that insufficient visual information from earlier stages propagates forward and induces object hallucination. Similarly, Kang et al. (2025) shows that only a small subset of attention heads, concentrated in the middle layers, are pivotal for visual grounding. Consistent with these findings, our analysis of visual representation alignment indicates that the preservation of visual information in the middle layers is strongly associated with spatial reasoning ability, which in turn is crucial for vision-centric tasks.
@@
While recent works have increasingly examined the internal information flow of MLLMs, most prior efforts remain concentrated on the input stage—particularly the use of frozen vision encoders. Improvements at this stage have largely focused on adopting stronger vision encoders (Kar et al., 2024; Lu et al., 2024; Shi et al., 2024; Azadani et al., 2025) or enhancing efficiency by reducing the overhead of visual tokens (Vasu et al., 2025; Yang et al., 2025; Wen et al., 2025). These advances have proven valuable, yet they primarily address the quality and efficiency of the initial visual representations, with comparatively less attention given to how visual information is processed and propagated once injected into the model. Recent efforts (Wang et al., 2024; 2025) take a step further by advocating direct supervision of visual tokens, but their focus remains on endpoint supervision with less consideration of the internal information flow. Moreover, their reconstruction-based objectives, while effective for preserving low-level fidelity, are less suited for capturing the higher-level semantic abstractions required by complex reasoning tasks (Zhang et al., 2023; Tong et al., 2024a).
@@ 3 Preliminaries
MLLMs typically consist of a pre-trained LLM ${LM}_{\theta}(\cdot)$ and a vision encoder $V_{\psi}(\cdot)$, which is connected with a vision-language projector $P_{\phi}(\cdot)$, where $\theta$, $\psi$, and $\phi$ denote corresponding learnable parameters. To generate answers grounded on both input image and text, the frozen vision encoder $V_{\psi}(\cdot)$ first extracts patch-level features from an input image $I\in\mathbb{R}^{H\times W\times 3}$ with height $H$ and width $W$ such that $\mathbf{z}=V_{\psi}(I)\in\mathbb{R}^{N\times D_{\mathbf{z}}}$, where $N$ and $D_{\mathbf{z}}$ denote the number of visual tokens and the dimension of the visual features, respectively. The subsequent projector $P_{\phi}(\cdot)$ maps these visual features into the language model’s embedding space, producing a sequence of visual tokens $\mathbf{e}^{\mathrm{img}}=P_{\phi}(\mathbf{z})\in\mathbb{R}^{N\times D}$, where $D$ denotes the hidden dimension of the language model. The text sequence is tokenized and embedded into the same embedding space using the language model’s token embedding layer, resulting in textual embeddings $\mathbf{e}^{\mathrm{text}}\in\mathbb{R}^{K\times D}$, where $K$ denotes the length of the text tokens. The language model then processes the concatenated multimodal sequence $\mathbf{e}^{\mathrm{img}};\mathbf{e}^{\mathrm{text}}]\in\mathbb{R}^{(N+K)\times D}$ and models the causal distribution over the text tokens $\mathbf{e}^{\mathrm{text}}$ as:
@@
To enable the language model to incorporate visual information, modern MLLMs typically follow a two-stage training paradigm (Liu et al., 2023; 2024): a vision–language pretraining stage followed by visual instruction tuning. Both stages share the same language-modeling objective but differ in parameter updates. During vision–language pretraining, only the projector parameters $\phi$ are optimized, while the language model parameters $\theta$ remain frozen. In contrast, visual instruction tuning jointly optimizes both $\phi$ and $\theta$, enabling the language model to adapt more deeply to visual inputs.
@@ 4.1 Do MLLMs undergo visual information loss?
While MLLMs ingest a substantial number of visual tokens, they are typically trained with a text-only language modeling loss applied to the output text tokens. Consequently, all learning signals are mediated through language supervision, and the visual representations $\mathbf{e}^{\mathrm{img}}$ receive no pathway-specific supervision, as illustrated in Fig. 2(a). In the absence of explicit visual supervision, we hypothesize that the model learns to prioritize only those visual features that immediately aid textual prediction, often discarding other potentially useful information. This, in turn, causes the internal visual representations to drift away from the rich features produced by the vision encoder—an effect that can undermine performance on tasks requiring complex reasoning or grounding.
To empirically validate this hypothesis, we measure the similarity between the internal visual representations of LLaVA (Liu et al., 2024) and the original visual features $\mathbf{z}$ extracted by its vision encoder (e.g., CLIP (Radford et al., 2021)). We adopt CKNNA (Huh et al., 2024) as a metric to quantify representational similarity.
As shown in Fig. 2(e), similarity to CLIP features drops sharply after the early layers and remains low in deeper layers, indicating that the model’s internal representations increasingly diverge from the encoder’s input features. This trend suggests that, without explicit visual supervision, the model has little incentive to preserve the encoder’s rich visual information.
Interestingly, despite the overall decline in alignment, we observe a modest recovery in the middle layers, suggesting that the network implicitly benefits from retaining visual representations at these depths when generating visually grounded answers. This observation aligns with prior analyses of information flow in MLLMs (Zhang et al., 2025b; Kaduri et al., 2025) and is also confirmed by our later layer-wise ablations, which show that leveraging the middle layers for vision-centric tasks shows the largest gains (see Section 2).
@@ 4.2 Is preserving visual information beneficial?
Figure 2: Re-injecting or aligning input visual features improves visual representation alignment and task performance. (a–d) Comparisons of (a) baseline visual instruction tuning (Liu et al., 2023), (b) re-injecting visual features from the post-projection layer, (c) re-injecting from the pre-projection layer, and (d) the proposed visual representation alignment, all applied at the 16-th layer. (e) Layer-wise alignment between visual tokens in MLLMs and vision encoder features measured by CKNNA (Huh et al., 2024), with shaded regions highlighting middle layers that are especially important for visual understanding. (f) Benchmark performance corresponding to (a–d).
Residual connection with post-projection features.
Having observed the mid-layer local increase in representation alignment, we ask whether explicitly preserving such visual information is beneficial. Let $\mathbf{e}^{\mathrm{img}}_{\ell}\in\mathbb{R}^{N\times D}$ denote the visual representations at the $\ell$-th layer of MLLMs. As a direct approach (Fig. 2(b)), we re-inject the projected visual representation $P_{\phi}(\mathbf{z})$ into an intermediate layer of the language model via a residual path:
$\mathbf{e}^{\mathrm{img}}_{\ell}\leftarrow\mathbf{e}^{\mathrm{img}}_{\ell}+P_{\phi}(\mathbf{z}).$ (3)
To isolate the effect of visual information retention without introducing new supervision, the model is trained solely with the original text loss $\mathcal{L}_{\mathrm{LM}}$. Unless otherwise stated, we set $\ell=16$ in a 32-layer model LLaVA (Liu et al., 2024), following our analysis that fine-grained visual understanding emerges most prominently in middle layers supported from later layer-wise ablations (see Section 2).
As shown in Fig. 2(e), the residual connection preserves alignment with the encoder’s visual features, as indicated by higher CKNNA similarity. Evaluated across standard multimodal benchmarks (Fig. 2(f)), this approach shows general improvements over the baseline, supporting the hypothesis that retaining encoder-aligned visual information benefits downstream tasks.
Although residual connection provides general gains, concerns remain that the vision-language projector, $P_{\phi}(\cdot)$, may not fully preserve the original visual information (Verma et al., 2024). This raises the question of whether using the encoder’s visual representations directly could better preserve visual information. To validate this hypothesis, we explore directions for connecting the raw encoder features directly to the language model in the following part.
Residual connection with pre-projection features.
We first explore connecting the input visual features to the language model via a residual branch using a lightweight adapter $A_{\phi^{\prime}}(\cdot)$ for dimensional compatibility. As illustrated in Fig. 2(c), we first conduct one such experiment by leveraging the vision encoder features $\mathbf{z}$ before the projector $P_{\phi}(\cdot)$ and re-injecting them into $\mathbf{e}^{\mathrm{img}}_{\ell}$ such that
$\mathbf{e}^{\mathrm{img}}_{\ell}\leftarrow\mathbf{e}^{\mathrm{img}}_{\ell}+A_{\phi^{\prime}}(\mathbf{z}).$ (4)
However, as shown in Fig. 2(f), this approach generally performs worse than the baseline. This is because the raw encoder features, which have not passed through the projector, are not sufficiently aligned with language features (Liu et al., 2023), and their direct residual connection consequently disrupts vision–language alignment in the intermediate layers. These findings suggest that incorporating external features into the internal visual pathway of LLMs requires more careful design.

### v2

@@ Abstract
Multimodal large language models (MLLMs) trained with visual instruction tuning have achieved strong performance across diverse tasks, yet they remain limited in vision-centric tasks such as object counting or spatial reasoning. We attribute this gap to the prevailing text-only supervision paradigm, which provides only indirect guidance for the visual pathway and often leads MLLMs to discard fine-grained visual details from the vision encoder during training. In this paper, we present VIsual Representation ALignment (VIRAL), a simple yet effective regularization strategy that aligns the internal visual representations of MLLMs with those of pre-trained vision foundation models (VFMs). By explicitly enforcing this alignment, VIRAL enables the model not only to retain critical visual details from its own vision encoder but also to complement additional visual knowledge from VFMs, thereby enhancing its ability to reason over complex visual inputs. Our experiments consistently demonstrate performance improvements across all tasks on widely adopted multimodal benchmarks, with gains reaching up to 17.3% and an average improvement of 9.4% over the baseline. Furthermore, we conduct comprehensive ablation studies to validate the key design choices underlying our framework. We believe this simple finding opens up an important direction for the effective integration of visual information in training MLLMs.
Figure 1: (a) VIsual Representation ALignment (VIRAL) introduces an auxiliary regularization objective on the visual pathway, preventing MLLMs from discarding detailed attributes of the input vision encoder during training while incorporating additional visual knowledge from vision foundation models (VFMs). (b) When trained with DINOv2 (Oquab et al., 2023) as the VFM, VIRAL consistently yields more accurate visually grounded responses and achieves substantial improvements over standard baselines (Liu et al., 2023) across diverse vision encoders, including CLIP (Radford et al., 2021) and SigLIPv2 (Tschannen et al., 2025).
@@ 1 Introduction
Recent advancements in multimodal large language models (MLLMs) (OpenAI, 2023; Bai et al., 2023a; Team et al., 2023; Chen et al., 2024d), particularly those employing visual instruction tuning techniques such as LLaVA (Liu et al., 2023), have achieved notable success in diverse multimodal tasks. By connecting pretrained large language models (LLMs) (Touvron et al., 2023; Chiang et al., 2023; Chen et al., 2024d; Bai et al., 2025) with vision encoders (Radford et al., 2021; Chen et al., 2024d; Tong et al., 2024a) through a lightweight vision–language projector, visual instruction tuning enables LLMs to interpret visual context and achieve strong performance across diverse tasks (Chen et al., 2024a; 2025a; Li et al., 2025a).
Despite these successes, numerous studies report persistent limitations in vision-centric tasks such as object counting and spatial reasoning (Tong et al., 2024b; Qi et al., 2025; Yuksekgonul et al., 2022; Ma et al., 2023). Early approaches largely attribute these shortcomings to the visual encoder or the projector. In response, subsequent works have introduced stronger vision encoders (Lu et al., 2024; Li et al., 2024) and more expressive projectors (Liu et al., 2024a; Cha et al., 2024; McKinzie et al., 2024), aiming to supply the language model with richer and more comprehensive visual representations. While they yield notable improvements, approaches that rely solely on more powerful vision encoders or projectors are inherently constrained in scalability and efficiency.
In this paper, we first revisit the conventional training paradigm of visual instruction tuning. Existing MLLMs are predominantly fine-tuned with a language-modeling objective, updating both the LLM and the vision-language projector while concentrating supervision almost entirely on textual outputs (Li et al., 2024; Bai et al., 2023b; Chen et al., 2024d). As a result, visual tokens receive only indirect, language-mediated supervision despite comprising a substantial fraction of the multimodal input. In effect, the visual pathway remains under-supervised, raising a central question: Is the prevailing multimodal training setup adequate for capturing and preserving visual information?
@@
Motivated by these findings, we propose VIsual Representation ALignment (VIRAL), a simple yet effective regularization strategy that directly supervises the visual pathway in MLLMs to prevent the model from discarding fine-grained visual attributes provided by the vision encoder during training. Specifically, we align the internal visual representations of the MLLMs with those of the initial vision encoder using an alignment loss based on cosine similarity. In addition, we further find that this alignment signal is much more effective when provided from stronger vision foundation models (VFMs) (Oquab et al., 2023; Kirillov et al., 2023; Yang et al., 2024; Ranzinger et al., 2024). Since VFMs are trained on vision-centric objectives, they provide rich visual representations that complement language supervision. Therefore, aligning the internal visual representations of MLLMs with those of VFMs likely allows the model to preserve critical visual details while also absorbing additional visual knowledge from VFMs, which in turn enhances its ability to reason over complex visual inputs. Through extensive experiments on widely adopted multimodal benchmarks, we show that VIRAL consistently delivers significant improvements across all tasks.
@@
  • We show that, under the visual instruction tuning paradigm, internal visual representations in MLLMs often lose alignment with the rich features produced by vision encoders, leading to the degradation of spatial reasoning capacity due to the loss of fine-grained visual information.
  • We propose a novel regularization strategy VIRAL, which explicitly aligns MLLM visual representations with features from pretrained VFMs, thereby preventing the loss of fine-grained attributes and enabling richer multimodal understanding.
  • Through comprehensive experiments on standard multimodal benchmarks, we show consistent and significant improvements of an average 9.4% over the baseline. In addition, we conduct extensive ablation studies and analysis to validate our design choices.
@@ 2 Related Work
Recent studies (Kaduri et al., 2025; Zhang et al., 2025b) have revealed a structured processing hierarchy in MLLMs for vision–language inputs: early layers aggregate global visual context into token embeddings, intermediate layers capture fine-grained spatial features, and later layers integrate multimodal information to facilitate response generation.
Within this hierarchy, the middle layers have been shown to be particularly critical for visual understanding. Jiang et al. (2025) decompose these layers into enrichment and refinement phases, showing that insufficient visual information from earlier stages propagates forward and induces object hallucination. Similarly, Kang et al. (2025) shows that only a small subset of attention heads in the middle layers are pivotal for visual grounding. Consistent with these findings, our analysis of visual representation alignment indicates that the preservation of visual information in the middle layers is strongly linked to spatial reasoning ability, which in turn is crucial for vision-centric tasks.
@@
While recent works have increasingly examined the internal information flow of MLLMs, most prior efforts remain concentrated on the input stage—particularly the use of frozen vision encoders. Improvements at this stage have largely focused on adopting stronger or multiple vision encoders (Kar et al., 2024; Lu et al., 2024; Shi et al., 2024; Azadani et al., 2025) or enhancing efficiency by reducing the overhead of visual tokens (Vasu et al., 2025; Yang et al., 2025; Wen et al., 2025). These advances have proven valuable, yet they primarily address the quality and efficiency of the initial visual representations, with comparatively less attention given to how visual information is processed and propagated once injected into the model. Recent efforts (Wang et al., 2024; 2025) take a step further by advocating direct supervision of visual tokens, but their focus remains on endpoint supervision with less consideration of the internal information flow. Moreover, their reconstruction-based objectives, while effective for preserving low-level fidelity, are less suited for capturing the higher-level semantic abstractions required by complex reasoning tasks (Zhang et al., 2023; Tong et al., 2024a).
@@ 3 Preliminaries
MLLMs typically consist of a pre-trained LLM ${LM}_{\theta}(\cdot)$ and a vision encoder $V_{\psi}(\cdot)$, which is connected with a vision-language projector $P_{\phi}(\cdot)$, where $\theta$, $\psi$, and $\phi$ denote corresponding learnable parameters. To generate answers grounded on both input image and text, the frozen vision encoder $V_{\psi}(\cdot)$ first extracts patch-level features from an input image $I\in\mathbb{R}^{H\times W\times 3}$ with height $H$ and width $W$ such that $\mathbf{z}=V_{\psi}(I)\in\mathbb{R}^{N\times D_{\mathbf{z}}}$, where $N$ and $D_{\mathbf{z}}$ denote the number of visual tokens and the dimension of the visual features, respectively. Projection modules vary across models—Resampler (Alayrac et al., 2022), Q-Former (Dai et al., 2023), and linear layers (Liu et al., 2023)—with linear layers recently dominating for their simplicity and strong performance. In this case, the linear projector $P_{\phi}(\cdot)$ maps these visual features into the language model’s embedding space, producing a sequence of visual tokens $\mathbf{e}^{\mathrm{img}}=P_{\phi}(\mathbf{z})\in\mathbb{R}^{N\times D}$, where $D$ denotes the hidden dimension of the language model. The text sequence is tokenized and embedded into the same embedding space using the language model’s token embedding layer, resulting in textual embeddings $\mathbf{e}^{\mathrm{text}}\in\mathbb{R}^{K\times D}$, where $K$ denotes the length of the text tokens. The language model then processes the concatenated multimodal sequence $\mathbf{e}^{\mathrm{img}};\mathbf{e}^{\mathrm{text}}]\in\mathbb{R}^{(N+K)\times D}$ and models the causal distribution over the text tokens $\mathbf{e}^{\mathrm{text}}$ as:
@@
To enable the language model to incorporate visual information, modern MLLMs typically follow a two-stage training paradigm (Liu et al., 2023; 2024a): a vision–language pretraining stage followed by visual instruction tuning. Both stages share the same language-modeling objective but differ in parameter updates. During vision–language pretraining, only the projector parameters $\phi$ are optimized, while the language model parameters $\theta$ remain frozen. In contrast, visual instruction tuning jointly optimizes both $\phi$ and $\theta$, enabling the language model to adapt more deeply to visual inputs.
@@ 4.1 Do MLLMs undergo visual information loss?
While MLLMs take a substantial number of visual tokens as input, they are typically trained with a text-only language modeling loss applied to the output text tokens. Consequently, all learning signals are mediated through language supervision, and the visual representations $\mathbf{e}^{\mathrm{img}}$ receive no vision-specific supervision, as illustrated in Figure 2-(a). In the absence of explicit visual supervision, we hypothesize that the model learns to prioritize only those visual features that immediately aid textual prediction, often discarding other potentially useful information. This, in turn, causes the internal visual representations to drift away from the rich features produced by the vision encoder—an effect that can undermine performance on tasks requiring complex visual reasoning or grounding.
To empirically validate this hypothesis, we measure the similarity between the internal visual representations of LLaVA (Liu et al., 2024a) and the original visual features $\mathbf{z}$ extracted by its vision encoder (e.g., CLIP (Radford et al., 2021)). We adopt CKNNA (Huh et al., 2024) as a metric to quantify representational similarity.
As shown in Figure 2-(d), similarity to CLIP features drops sharply after the early layers and remains low in deeper layers, indicating that the model’s internal visual representations increasingly diverge from the encoder’s input features. This trend suggests that, without explicit visual supervision, the model has little incentive to preserve the encoder’s rich visual information.
Interestingly, despite the overall decline in alignment, the middle layers show a clear attenuation of this trend, with even slight increase, suggesting that the network implicitly benefits from retaining visual representations at these depths when generating visually grounded answers. This observation aligns with prior analyses of information flow in MLLMs (Zhang et al., 2025b; Kaduri et al., 2025) and is also confirmed by our later layer-wise ablations, which show that leveraging the middle layers for vision-centric tasks shows the largest gains (see Section 5.3).
Figure 2: Re-injecting or aligning visual features improves representation alignment and performance. (a–c) Comparison of (a) baseline visual instruction tuning (Liu et al., 2023), (b) re-injecting visual features, and (c) visual representation alignment, all applied at the 16th layer. (d) Layer-wise alignment between visual tokens in MLLMs and vision encoder features, measured by CKNNA (Huh et al., 2024), with shaded regions denoting middle layers that are particularly important for visual understanding. (e) Benchmark performance corresponding to (a–c).
@@ 4.2 Is preserving visual information beneficial?
Having observed the mid-layer local increase in representation alignment, we ask whether explicitly preserving such visual information is beneficial. Let $\mathbf{e}^{\mathrm{img}}_{\ell}\in\mathbb{R}^{N\times D}$ denote the visual representations at the $\ell$-th layer of MLLMs. As a direct approach (Figure 2-(b)), we re-inject the projected visual representation $P_{\phi}(\mathbf{z})$ into an intermediate layer of the language model via a residual path:
$\mathbf{e}^{\mathrm{img}}_{\ell,i}\leftarrow\mathbf{e}^{\mathrm{img}}_{\ell,i}+P_{\phi}(\mathbf{z}_{i}).$ (3)
To isolate the effect of visual information retention without introducing new supervision, the model is trained solely with the original text loss $\mathcal{L}_{\mathrm{LM}}$. Unless otherwise stated, we set $\ell=16$ in a 32-layer model LLaVA (Liu et al., 2024a), following our analysis that fine-grained visual understanding emerges most prominently in middle layers, supported by later layer-wise ablations (see Section 5.3).
As shown in Figure 2-(d), adding the residual connection better preserves the alignment with the encoder’s visual features, as indicated by higher CKNNA similarity. Evaluated across standard benchmarks (Figure 2-(e)), this approach shows general improvements over the baseline, supporting the hypothesis that retaining encoder-aligned visual information benefits downstream tasks.
Although residual connection provides general gains, concerns remain that the vision-language projector, $P_{\phi}(\cdot)$, may not fully preserve the original visual information (Verma et al., 2024; Cha et al., 2024). This raises the question of whether using the encoder’s visual representations directly could better preserve visual information. To validate this hypothesis, we explore directions for connecting the raw encoder features directly to the language model in the following part.